Find the percentage decrease in area of a triangle if it's each side is halved
 

Grade:9

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
Assuming the triangle to be a rigth angled triangle and sides to be AB,BC and CA and let angle
then previous area= ½ * AB * BC
 
Now all sides are halved so new triangle is A’B’C’ with A’B’=AB/2, B’C’=BC/2 and C’A’=CA/2.
New area = ½ * A’B’ * B’C’
=1/2*(AB/2)*(BC/2)
=1/2*1/2*previous area
=1/4 * previous area
 
So decrease = previous area-new area = 3/4*previous area
 
percentage decrease in area = ¾ * 100 = 75%
 
So the percentage decrease in area is 75%.
 
Hope this solved your question.
Thank you,
Soumya Ranjan Mohanty(IIIT Bhubaneswar).